# The role of spectator modes in the quantum-logic spectroscopy of single trapped molecular ions

### v1

@@ Abstract
Quantum-logic spectroscopy has become an increasingly important tool for the state detection and readout of trapped atomic and molecular ions which do not possess easily accessible closed-optical-cycling transitions. In this approach, the internal state of the target ion is mapped onto a co-trapped auxiliary ion. This mapping is typically mediated by normal modes of motion of the two-ion Coulomb crystal in the trap. The present study investigates the role of spectator modes not directly involved in a measurement protocol relying on a state-dependent optical-dipole force. We identify a Debye-Waller-type effect that modifies the response of the two-ion string to the force and show that cooling all normal modes of the string allows for the detection of the rovibrational ground state of a N molecular ion with a fidelity exceeding 99.99% improving on previous experiments by more than an order of magnitude. This marked improvement in sensitivity paves the way for simultaneously identifying multiple rovibrational states at a fixed set of experimental parameters.
@@
Quantum-logic spectroscopy (QLS) protocols provide a framework for the readout of the internal state of a target (‘spectroscopy’) ion by mapping it onto the state of a co-trapped axillary (‘logic’) ion [2]. This technique is usually employed for systems without readily accessible optical cycling transitions which are usually used for state detection, e.g., non-laser-coolable atomic species such as Al+ [2], highly-charged ions [3] and molecular ions [4, 5, 6, 7]. The state mapping is typically mediated by a motional degree of freedom common to the ions which are strongly coupled by the Coulomb interaction. In general, such a QLS experimental sequence consists of three steps: first, cooling of a common (target) motional mode, second, manipulation of the target mode population depending on the internal state of the spectroscopy ion and third, motional-state readout by spectroscopy on the logic ion. The QLS implementations demonstrated so far used different approaches to project the state of the spectroscopy ion onto the motional state. Several experiments involved adding a phonon to a target mode initially cooled to its ground state by driving a transition on a motional sideband of a spectroscopic transition in the spectroscopy ion [8, 5, 3]. In such a scheme, the addition of a phonon to the target mode indicated a successful state detection. In another approach, the target mode was prepared in the first excited motional Fock state and, depending on the internal state of the spectroscopy ion, a single phonon was exchanged with another mode [4]. This mode transfer was mediated by a state-dependent optical dipole force (ODF) resonantly driving the mode exchange. Both of these approaches involved single-phonon excitations. A third method relies on a modulated state-dependent ODF tuned to resonance with the target motional mode, thus causing coherent motional excitation [9, 6]. Depending on the duration of the ODF pulse, the target mode can thus be excited to highly excited motional states. Here, we follow this latter approach. Various alternative implementations of QLS were theoretically proposed, but remain to be demonstrated [10, 11, 12, 13].
In the present paper, we further characterize a quantum-logic spectroscopy (QLS) protocol used to determine the internal state of single N molecular ions (the spectroscopy ion) [14, 6, 15]. A co-trapped Ca+ ion serves as a coolant and a probe. We use a running optical lattice comprised of two counter-propagating laser beams around 787 nm near-resonant with specific rovibronic transitions in order to realize a state-dependent ODF on the N ion. The optical lattice is modulated at a frequency resonant with a normal mode of the two-ion string. The resulting resonant motional excitation of the ions in this mode is probed with sideband Rabi spectroscopy on the logic ion [16, 17] and thus provides information about the state of the spectroscopy ion.
The motion of a two-ion crystal in a linear radiofrequency (RF) ion trap (see schematic in Fig. 1) can be described in terms of six normal modes. Two modes, with in-phase (ax-IP) and out-of-phase (ax-OP) motions of the ions, are directed along the longitudinal trap axis. The other four modes are directed along two perpendicular radial principal axes of the trap (rad-OP and rad-IP along each axis). Due to small asymmetries of the radial potential, the corresponding radial mode frequencies usually differ slightly ( 10 kHz here).
@@
Cooling the target mode close to its quantum-mechanical ground state (here, we use the axial centre-of-mass mode, ax-IP) is a prerequisite for the present implementation of QLS mediated with motional states. However, population in the remaining normal modes which do not directly participate in the measurement protocol (spectator modes) can modify the light-ion interactions via the Debye-Waller (DW) effect [18]. The DW effect manifests itself in a change of Rabi frequencies of transitions and introduces shot-to-shot frequency fluctuations and shifts which are of relevance for applications in precision spectroscopy, quantum simulations, and quantum computing with multiple ions [18, 2, 19, 20].
@@
DW effects are present in two stages of the present QLS method. First, population in the spectator modes may alter the interaction of the molecule with the ODF pulse. Second, it also affects the readout of the motional excitation with sideband spectroscopy leading to a reduced signal-to-noise ratio and, thus, decreased state-detection fidelities. While the second effect was shown to be small in previous studies [6], the first can have significant ramifications for the state-detection fidelity, as shown here.
@@
Here, differences in response to the ODF pulse when the two-ion Coulomb crystal was prepared in specified spectator mode states were observed experimentally and rationalized with the aid of simulations. It is shown that the spectator modes play an important role during the motional excitation via the DW effect. By cooling all spectator modes close to their ground states, we achieve an improved state-detection fidelity, exceeding 99.99% for as few as eight experimental repetitions, compared to 99.5% after 22 cycles reported previously when only the target mode was cooled [6]. Furthermore, we show that the improved sensitivity of the detection method can also assist in the identification of the higher-lying rotational states.
2 Experimental Methods
@@
Single 14N ions were prepared in the rovibrational ground state using a 2+1′ resonance enhanced multi-photon ionization (REMPI) scheme of internally cold neutral N2 molecules from a supersonic molecular beam [21]. The molecular ions were loaded into a linear trap operated at an RF of 19.4 MHz and endcap voltages of 150 V. The motional frequencies for the normal modes of the N-Ca+ ion string, , were measured using spectroscopy of motional sidebands on the Ca+ Zeeman-resolved ‘clock’ transition around 729 nm and are presented in Table 1. A static external magnetic field of 4.70 G was applied to define the quantization axis.
@@
With the present laser configuration, we were not able to resolve the rad-IP modes due to the small power of the interrogation laser (729 nm) in the radial direction and the current transition coherence times in the setup. Therefore, the average phonon number after cooling for the rad-IP modes quoted in Table 1 is an estimate based on (1) the strongly increased -time for sideband transitions on this mode after EIT cooling and (2) the estimated performance of the EIT cooling for this mode (see sec. 2.1 below).
@@ 2.1 Cooling methods
Schematics of the geometric configuration of the lasers with respect to the ion trap as well as the energy levels and transitions used for the present cooling schemes are depicted in Figs. 1 (a) and (b), respectively. Laser cooling was performed on Ca+ in two stages. First, all six modes of the two-ion crystal were cooled close to the Doppler limit by driving a red-detuned transition with a 397 nm laser beam with a 45∘ projection along the axial and 60∘ projection to the radial axes. Another laser at 866 nm served to repump population from the metastable D3/2 level to close the cooling cycle. Next, the ions were prepared close to the motional ground state of the target (ax-IP) mode using different secondary cooling schemes:
@@
Pulsed sideband cooling (PSC) utilises a train of excitation pulses on the second and first-order sidebands on a narrow transition [22, 17], here ax-IP red sidebands (RSB) on the Ca+ clock transition. A repumper laser pulse at 854 nm followed each of the sideband pulses to pump the ion from the metastable state to the state from where it decayed back to the electronic ground state (Fig. 1 (b)). The PSC sequence is shown in Fig. 1 (c). The number and duration of each pulse were empirically optimised to provide the best performance. Moreover, if required, either one (SB) or two (SB) additional cooling pulse sequences on the ax-OP (spectator) mode were interlaced in the SB cooling sequence to reduce the population in this mode down to three or below one phonon. If no pulses were added (SB) the average ax-OP mode population was close to the Doppler limit.
@@
EIT cooling [23, 24, 25] was realized on Zeeman components of the 397 nm transition between two and upper levels with a -polarized coupling beam and two counter-propagating -polarized cooling beams (Fig. 1 (b)). Each of the cooling laser beams provided a projection of its -vector along one of the radial directions and at 45∘ to the axial direction. In this configuration, the axial modes were always cooled. However, by using only one or both EIT -polarized cooling beams, either one (EIT2) or both radial modes (EIT3) were additionally targeted.
@@
In these experiments, the motional modes (see Table 1) could be divided into two frequency groups – first, around 600 kHz, with ax-IP and rad-OP modes, and second, around twice the frequency of the first group with ax-OP and rad-IP. Two EIT pulses with adjusted detuning of the involved lasers as well as the intensity of the -polarized coupling laser beam were implemented to cool each of the groups efficiently [25, 26]. In Fig. 1 (d), an example of such a multi-stage EIT cooling sequence is presented. Each pulse is annotated with the axial mode around which the EIT resonance was placed, ensuring optimal cooling.
@@
It was experimentally confirmed that, given sufficient EIT cooling time, all modes could be efficiently cooled below one phonon. All three EIT lasers were detuned from the state by 100 MHz in the first stage and then by 80 MHz in the second one. EIT cooling reduced the population of the ax-IP mode to 0.4 phonons. To cool this mode even further, the EIT sequence was followed by a set of first-RSB pulses similar to the ones used in PSC to ensure a similar minimum temperature of the ax-IP mode across all cooling methods.
@@ 2.2 Motional mode populations after cooling
The cooling speed is an interplay of the cooling and heating processes which are dominated by stray electric fields and electric noise. The heating was measured to be 0.05 phonons/ms for the ax-IP and 0.1 phonons/ms for the ax-OP mode [30]. These heating rates are negligible compared to the timescales of the present experiments.
@@ 2.3 Motional excitation
The ODF generating the motional excitation of the two-ion string was realized using a running lattice generated by two counter-propagating -polarized laser beams at 787.4505 nm referenced to a wavemeter (HighFinesse WSU-30). The wavelength of the lattice lasers was detuned by 12 GHz from the rovibronic transition in N. Because the ac-Stark effect causing the ODF is inversely proportional to the detuning from the transition, the lattice interacts strongest with, and is thus most sensitive to, a molecule in the ground rovibronic state [6, 31]. Before applying the lattice lasers, the Ca+ ion was shelved in the level to suppress spurious ODF generated by the atomic ion [6]. The ODF pulse length was set to 500 s for all experiments presented here.
@@ 2.4 State detection
The state-dependent ODF projected the internal molecular state on the ax-IP motional state of the two-ion crystal, which was subsequently probed by Rabi blue-sideband spectroscopy on the Ca+ clock transition [6]. A typical blue-sideband Rabi flop is shown in Fig. 2. If the ax-IP mode was previously cooled to its motional ground state, a Rabi flop was only observed when motional excitation occurred, i.e., when the molecule was in the rovibronic ground state (orange trace). When N was in an excited state, there were no oscillations (violet trace), similar to the background signal (blue).
As the probe laser was aligned along the trap axis, only the axial modes contribute directly to the signal. The target (ax-IP) mode population, , and the spectator (ax-OP) mode population, , affect the Rabi frequency for a blue-sideband transition in a two-level system according to [18, 14]:
where is the bare Rabi frequency and a generalised Lauguerre polynomial. The Lamb-Dicke parameter for the target mode () is , and for the spectator mode () is . The spatial extent of the motional-ground-state wavefunction for Ca+, for each mode is:
The frequencies of the two normal modes are given by [32, 33]:
with the axial frequency of a single Ca+ ion of mass in the trap and the mass ratio . The mass-dependent factor is defined as:
The populations in the spectator modes introduce corrections to the Rabi frequency in the form of a DW factor [18]:
The ground-state population for a two-level system undergoing Rabi oscillations is weighted over the probabilities of the ion occupying different motional states, , and can be written as [18]:
where is the pulse time and is the detuning of the excitation laser frequency (in the present experiments, ). The depend on the applied cooling methods, on heating effects and on the characteristics of the motional excitation.
Additionally, the Rabi flop is subject to decoherence. Unlike the lifetime of the excited metastable state of Ca+ that is of the order of a second, the decoherence time for a motional sideband transition when trapping two-ion string, is around 500 s and has to be included in the fitting of experimental Rabi flops. To account for this, and for the fact that in the presented experiments Ca+ was initially shelved in , the excited-state probability used for fitting the experimental data is [18]:
@@ 3 Simulations
To interpret the experimental results, we performed both classical and quantum simulations of the two-ion motional-excitation dynamics when applying the ODF with the running lattice. The simulations yield motional population distributions, represented as from Eq. (6), for different lattice parameters. Here, we briefly discuss the theory underlying the simulations. Although solutions for two ions are presented for generality, we emphasize that in the experiments and simulations, only N interacted with the lattice.
@@
Since the radial modes perpendicular to the axial direction of propagation of the running lattice are not expected to contribute to the excitation dynamics (as confirmed experimentally, see below), the simulations are restricted to the axial direction.
@@
The classical one-dimensional equation of motion for each ion of mass interacting with the running optical lattice pulse lasting is given by [14]:
where is the frequency of ion in the trap, is the wavenumber of a single lattice laser, is the frequency difference between the lattice lasers, is the elementary charge, is the vacuum permittivity, and are the -coordinates of the ions. is the ac-Stark shift on ion exerted by a single-lattice laser beam. The terms in the equation correspond to the interaction of the ion with the trapping potential, with the co-trapped ion, and with the optical lattice, respectively.
@@
Note the factor of two which originates from using two counter-propagating lasers that make up the running lattice. The DW factors in Eq. (3) are defined as:
To obtain and hence calculate the mode populations , the Schrödinger equation can be solved either directly numerically starting from the Hamiltonian Eq. (9) and using, e.g., the Qutip package [36], or by solving the system of differential equations Eq. (3). The latter method is significantly less expensive computationally. However, it assumes only single-phonon transitions in the target mode, no spectator-mode excitation during the dynamics and no cross-talk between the modes. These conditions are satisfied when the detuning of the optical lattice from the target-mode frequency is smaller than the light frequency, the detuning from the spectator mode and both normal mode frequencies, i.e. .
Such an assumption may break down for high average phonon populations and broader phonon distributions when the probability of exciting non-linear resonances may become significant [37]. Hence, we tested the equivalence of the two simulation methods for a relevant range of excitation times and ac-Stark shifts and observed no significant discrepancy between the models. Hence, we used the faster method capitalizing on Eq. (3) to obtain the motional state distributions required for simulating the experimental blue-sideband Rabi oscillations Eq. (7). In the simulations, the ions were initially in the ground state of the target mode, with different populations on the OP spectator mode, i.e. .

### v2

@@ Abstract
Quantum-logic spectroscopy has become an increasingly important tool for the state detection and readout of trapped atomic and molecular ions which do not possess easily accessible closed-cycling optical transitions. In this approach, the internal state of the target ion is mapped onto a co-trapped auxiliary ion. This mapping is typically mediated by normal modes of motion of the two-ion Coulomb crystal in the trap. The present study investigates the role of spectator modes not directly involved in a measurement protocol relying on a state-dependent optical-dipole force. We identify a Debye-Waller-type effect that modifies the response of the two-ion string to the force. We show that cooling the spectator modes of the string allows for the detection of the rovibrational ground state of an N molecular ion with a computed statistical fidelity exceeding 99.99%, improving on previous experiments by more than an order of magnitude while also halving the experimental time. This enhanced sensitivity enables the simultaneous identification of multiple rotational states with markedly weaker signals.
@@
Quantum-logic spectroscopy (QLS) protocols provide a framework for the readout of the internal state of a target (‘spectroscopy’) ion by mapping it onto the state of a co-trapped axillary (‘logic’) ion [1]. This technique is usually employed for systems without readily accessible optical cycling transitions, e.g., atomic species such as Al+ [1], highly-charged ions [2] and molecular ions [3, 4, 5, 6]. The state mapping is typically mediated by a motional degree of freedom common to the ions which are strongly coupled by the Coulomb interaction. In general, a QLS experimental sequence consists of three steps: first, cooling of a common (target) motional mode, second, manipulation of the target mode population depending on the internal state of the spectroscopy ion and third, motional-state readout by spectroscopy on the logic ion.
The QLS implementations demonstrated so far used different approaches to project the state of the spectroscopy ion onto the motional state. Several experiments involved adding a phonon to a target mode initially cooled to its ground state by driving a transition on a motional sideband of a spectroscopic transition in the spectroscopy ion [7, 6, 4, 2]. In such a scheme, the addition of a phonon to the target mode indicated a successful state detection. In another approach, the target mode was prepared in the first excited motional Fock state, , and depending on the internal state of the spectroscopy ion, a single phonon was exchanged with another mode [3]. This mode transfer was mediated by a state-dependent optical dipole force (ODF) resonantly driving the mode exchange. Both of these approaches involved single-phonon excitations. A third method relies on a modulated state-dependent ODF tuned to resonance with the target motional mode, thus causing motional excitation [8, 5]. Depending on the duration of the ODF pulse and the state of the spectroscopy ion, the target mode can thus be excited to highly excited motional states. Various alternative implementations of QLS were proposed [9, 10, 11, 12].
Here, we further explore the third QLS method mentioned above, which employs a state-dependent ODF to determine the internal state of the spectroscopy ion [8, 13, 5, 14]. The ODF is modulated at a frequency resonant with a normal mode of the two-ion system (referred to as target mode) and its amplitude depends on the internal state of the spectroscopy ion. The resulting resonant motional excitation of the system’s motional mode is probed by the logic ion using sideband spectroscopy [15, 16] and thus provides information about the state of the spectroscopy ion.
@@
Cooling the target mode close to its quantum-mechanical ground state is a prerequisite for most implementations of QLS mediated with motional states. However, population in the remaining normal modes which do not directly participate in the measurement protocol (spectator modes) could have detrimental effects on the spectroscopy due to Kerr-like [17], mode-cross-talk [18, 19] or Debye-Waller (DW) [18] effects. Consequently, it is common to set the frequencies of all the modes to be well-separated (i.e. avoid any resonances), and to cool all modes to sub-Doppler temperatures [4, 20]. Here, we study in detail the consequences for QLS of the DW effect that modifies light-ion interaction strength due to motion of the ion. This effect was shown to be of relevance for applications in precision spectroscopy, quantum simulations, and quantum computing with multiple ions [18, 1, 21, 22, 23, 24].
@@
DW effects are present in two stages of the present QLS method. First, population in the spectator modes may alter the interaction of the molecule with the ODF pulse. Second, it also affects the readout of the motional excitation with sideband spectroscopy, leading to a reduced signal-to-noise ratio and, thus, decreased state-detection fidelities. While the second effect was shown to be small in previous studies [5], the first can have significant ramifications for the state-detection fidelity, as shown here.
@@
We experimentally observe differences in response to the ODF pulse when the two-ion Coulomb crystal was prepared in specified spectator mode states. We rationalized these differences with the aid of analytical models and numerical simulations. We show that the spectator modes play an important role during the motional excitation via the DW effect. By cooling the axial spectator mode close to its ground state, we achieve an improved computed statistical state-detection fidelity, exceeding 99.99% for as few as nine experimental repetitions. Compared to 99.5% after 22 cycles reported previously when only the target mode was cooled [5], the execution time of the protocol was halved, while simultaneously achieving a more precise state determination. Furthermore, we show that the improved sensitivity of the detection method can also assist in the identification of higher-lying rotational states of the molecule.
2 Methods
@@
Single 14N ions were prepared in the rovibrational ground state using a 2+1′ resonance enhanced multi-photon ionization (REMPI) scheme of internally cold neutral N2 molecules from a supersonic molecular beam [25]. The molecular ions were loaded into a linear trap operated at a radiofrequency of 19.1 MHz and endcap voltages of 150 V. The motional frequencies for the normal modes of the N-Ca+ ion string, , were measured using spectroscopy of motional sidebands on the Ca+ Zeeman-resolved ‘clock’ transition around 729 nm and are presented in Table 1. A static external magnetic field of 4.70 G was applied to define the quantization axis.
The motion of a two-ion crystal in a linear radiofrequency ion trap (see schematic in Fig. 1) can be described in terms of six normal modes. Two modes, with in-phase (ax-IP) and out-of-phase (ax-OP) motions of the ions, are directed along the longitudinal trap axis. The other four modes are directed along two perpendicular radial principal axes of the trap (rad-OP and rad-IP along each axis). Due to small asymmetries of the radial potential, the corresponding radial mode frequencies usually differ slightly ( 10 kHz here).
@@ 2.1 Cooling methods
Schematics of the geometric configuration of the lasers with respect to the ion trap as well as the energy levels and transitions used for the present cooling schemes are depicted in Figs. 1 (a) and (b), respectively. Laser cooling was performed on Ca+ in two stages. First, all six modes of the two-ion crystal were cooled close to the Doppler limit by driving a red-detuned transition with a 397 nm laser beam with a 45∘ projection along the axial and 60∘ projection to the radial axes. Another laser at 866 nm served to repump population from the metastable level, closing the cooling cycle. Next, the ions were prepared close to the motional ground state of the target (ax-IP) mode using different secondary cooling schemes:
@@
Pulsed sideband cooling (PSC) utilises a train of excitation pulses on the second and first-order sidebands on a narrow transition [16], here ax-IP red sidebands (RSB) on the Ca+ clock transition. A repumper laser pulse at 854 nm followed each of the sideband pulses to pump the ion from the metastable state to the state, from where it decayed back to the electronic ground state (Fig. 1 (b)). The PSC sequence is shown in Fig. 1 (c). The number and duration of each pulse were empirically optimised to provide the best cooling performance. Moreover, if required, either one (SB) or two (SB) additional cooling pulse sequences on the ax-OP (spectator) mode were interlaced in the SB cooling sequence to reduce the population in this mode down to three or below one phonon. If no pulses were added (SB), the average ax-OP mode population was close to the Doppler limit.
@@
EIT cooling [26, 27, 28] was realized on Zeeman components of the 397 nm transition between the and levels with a -polarized coupling beam and two counter-propagating -polarized cooling beams (Fig. 1 (b)). Each of the cooling laser beams provided a strong projection of its -vector along one of the radial directions and at 60∘ to the axial direction. In this configuration, the axial modes were always cooled. However, by using only one or both EIT -polarized cooling beams, modes along either one (EIT2) or both radial axes (EIT3) were efficiently addressed.
@@
In our experiments, the motional modes (see Table 1) could be divided into two frequency groups – first, around 600 kHz, with ax-IP and rad-OP modes, and second, around twice the frequency of the first group with ax-OP and rad-IP. Typically, two EIT pulses were implemented to cool each of the groups efficiently [28, 29]. In Fig. 1 (d), an example of such a multi-stage EIT cooling sequence is presented. Each pulse is annotated with the axial mode targeted for optimal cooling.
@@
It was experimentally confirmed that, given sufficient EIT cooling time, axial and rad-OP modes could be efficiently cooled below one phonon. The EIT parameters for each stage were adjusted empirically to provide optimal cooling for the modes of interest. EIT cooling reduced the population of the ax-IP mode to 0.4 phonons. To cool this mode even further, the EIT sequence was followed by a set of first-RSB pulses similar to the ones used in PSC to ensure a similar minimum temperature of the ax-IP mode across all cooling methods.
@@ 2.2 Motional mode populations after cooling
The heating rates in our trap were measured to be 0.05 phonons/ms for the ax-IP and 0.1 phonons/ms for the ax-OP mode [33]. These heating rates are negligible compared to the timescales of the present experiments. The temperature and heating rates of the radial modes were observed to have a negligible effect on the axial excitation dynamics due to their negligible Lamb-Dicke factors with the lattice lasers which were generally directed along the longitudinal trap axis.
Generally, the 729 nm interrogation laser beam was aligned along the longitudinal trap axis thus only addressing the axial modes. For diagnostics of the radial modes, we aligned part of the laser beam to a 60∘ projection onto the radial modes (see Fig. 1 (a)). The average phonon number after cooling for the rad-IP modes quoted in Table 1 is an estimate based on the expected performance of the EIT cooling for this mode.
@@ 2.3 Motional excitation
The ODF generating the motional excitation of the two-ion string was realized using a travelling lattice generated by two counter-propagating -polarized laser beams with a frequency difference of the IP axial mode between them. The wavelength of the lattice laser (787.4505 nm referenced to a HighFinesse WSU-30 wave meter) was detuned by 12 GHz from the rovibronic transition in N. Because the ac-Stark effect causing the ODF is inversely proportional to the detuning from the transition, the lattice interacts strongest with, and is thus most sensitive to, a molecule in the ground rovibronic state [5, 34]. Before applying the lattice lasers, the Ca+ ion was shelved in the level to suppress spurious ODF generated by the atomic ion coupling to the states [5]. The ODF pulse length was set to 500 s for all experiments presented here, except of the example in Fig. 2 (a) for which it was 1 ms and the experiments in Fig. 7 for which it was 2 ms.
@@ 2.4 State detection
The state-dependent ODF projected the internal molecular state on the ax-IP motional state of the two-ion crystal, which was subsequently probed by Rabi blue-sideband spectroscopy on the Ca+ clock transition [5]. An example blue-sideband Rabi flop is shown in Fig. 2 (a). When the ax-IP mode is initially cooled to its motional ground state, a Rabi flop is observed only when motional excitation occurs, i.e., when the molecule is in the rovibronic ground state (orange trace). When N is in an excited state, there are no oscillations (violet trace), similar to the background signal (blue).
As the probe laser was aligned along the trap axis, only the axial modes contribute directly to the signal. The target (ax-IP) mode population, , and the spectator (ax-OP) mode population, , affect the Rabi frequency for a blue-sideband transition in a two-level system [18, 13]:
where is the bare Rabi frequency and a generalised Lauguerre polynomial. Note that since the Ca+ is initialized in the excited electronic state, the blue sideband corresponds to a change in the target mode of . The Lamb-Dicke parameter for the target mode () is , and for the spectator mode () is , taking values 0.096 and 0.051 respectively. The spatial extent of the motional-ground-state wavefunction for Ca+ for each mode is
where the frequencies of the two normal modes are given by [35, 36]:
with the axial frequency of a single Ca+ ion of mass and the mass ratio of the ions is given by . The mass-dependent factor is defined as:
In Eq. 1, the populations in the spectator modes introduce corrections to the Rabi frequency in the form of a DW factor [18]:
The excited-state population for a two-level system undergoing Rabi oscillations is weighted over the probabilities of the ion occupying different motional states, , and can be written as [18]:
where is the pulse time and is the detuning of the excitation laser frequency from the blue-sideband transition (in the present experiments, ). The probabilities depend on the applied cooling methods, on heating effects, and on the characteristics of the motional excitation, as elaborated in Sec. 3.
Additionally, the Rabi flop is subject to decoherence. Unlike the lifetime of the excited metastable state of Ca+ that is of the order of a second, the decoherence time for the motional sideband transition when trapping a two-ion string, , is around 500 s and has to be included in the fitting of experimental Rabi flops. To account for this, and for the fact that in the presented experiments Ca+ was initially shelved in , the excited-state probability used for fitting the experimental data is [18]:
@@ 3 Simulations
To interpret the experimental results, we performed both classical and quantum simulations of the two-ion motional-excitation dynamics when applying the ODF with the travelling lattice. The simulations yield motional population distributions, represented as from Eq. (6), for different lattice parameters. Here, we briefly discuss the theory underlying the simulations. Although solutions for two ions are presented for generality, we emphasize that in the experiments and simulations, only N interacted with the lattice.
@@
Since the radial modes are perpendicular to the axial direction of propagation of the travelling lattice, they are not expected to contribute to the excitation dynamics (as confirmed experimentally, see Fig. 4 (b)). Hence, the simulations are restricted to the axial direction.
@@
The classical one-dimensional equation of motion for each ion of mass interacting with the travelling-optical-lattice pulse is given by [13]:
where is the frequency of ion in the trap, is the wavenumber of a single lattice laser, is the frequency difference between the lattice lasers, is the elementary charge, is the vacuum permittivity, and are the -coordinates of the ions. is the ac-Stark shift on ion exerted by a single lattice-laser beam. The terms in the equation correspond to the interaction of the ion with the trapping potential, with the co-trapped ion, and with the optical lattice, respectively. The sign before the second term is ‘-’ for =1 and ‘+’ for =2 respectively.
@@
Note the factor of two which originates from using two counter-propagating lasers forming the travelling lattice. The mass-dependent factor is defined according to Eq. (4). The Lamb-Dicke parameters thus defined for the presented experiments took values = (0.160, 0.150, -0.094, 0.179) for the presented experiments.
The DW factors of the spectator mode in Eq. (3) are defined as:
From this definition it becomes apparent that the DW factors act as damping factors for the motional excitation which depend on the spectator-mode populations. In Sec. 4, we will see their significant effect on the motional Fock-state distribution after applying the ODF.
To obtain and hence calculate the mode populations, , the Schrödinger equation can be solved either directly numerically starting from the Hamiltonian Eq. (9) and using, e.g., the Qutip package [39], or by solving the system of differential equations Eq. (3). The latter method is significantly less computationally expensive. However, it assumes only single-phonon transitions in the target mode, no spectator-mode excitation during the dynamics and no cross-talk between the modes. These conditions are satisfied when the detuning of the optical lattice from the target-mode frequency is smaller than the detuning from the spectator mode and both normal mode frequencies, i.e. .
These assumptions may break down for high average phonon populations and broader phonon distributions when the probability of exciting non-linear resonances may become significant [40]. Hence, we tested the equivalence of the two simulation methods for a relevant range of excitation times and ac-Stark shifts and observed no significant discrepancy between the models. Hence, we used the faster method capitalizing on Eq. (3) to obtain the motional state distributions required for simulating the experimental blue-sideband Rabi oscillations Eq. (7). In the simulations, the ions were initially in the ground state of the target mode with different populations on the OP spectator mode.
In the considered model, we do not include non-linear effects originating from the higher-order terms in the Coulomb interaction between the ions [17, 41]. Such “Kerr-like” effects are known to modulate ax-OP frequencies by coupling to rad-OP modes (or rad-OP modes to each other) in the case of equal-mass two-ion strings [42]. Also, if these modes fall into parametric resonance conditions, such as , the interaction anharmonicities can lead to a non-negligible population transfer.